I'm trying to help my son with some highschool algebra. I am wondering if you can explain two problems
2x=4y
6+y=8
The question is determine whether each equation is a linear equation, if so write it in standaard form. Please help

Answer by checkley75(3666)

2X=4Y
4Y=2X
Y=2X/4
Y=X/2
-----------------------
6+Y=8
Y=8-6
Y=2
-----------------------
THESE TWO ARE LINEAR EQUATIONS.